# Data-driven Model Predictive Control: Asymptotic Stability despite Approximation Errors exemplified in the Koopman framework

### v2

@@ 3. Asymptotic stability of MPC
Since, for each $i\in[0:N-1]$, we have the inequality
$\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+}),u^{\sharp}(i))\geq\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|_{Q}^{2},$
we also have $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|_{Q}^{2}\leq B^{\varepsilon}_{N}\|\hat{x}\|_{Q}^{2}$. By exploiting standard inequalities of weighted squared norms and taking the square root, we have $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|\leq\sqrt{B^{\varepsilon}_{N}\bar{\lambda}/\underline{\lambda}}\|\hat{x}\|$. Then, substituting this inequalities in (15), we get
@@
$\leq$ $)\displaystyle 2\|Q\|\sqrt{B^{\varepsilon}_{N}\frac{\bar{\lambda}}{\underline{\lambda}}}\|\hat{x}\|L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}\|\hat{x}\|+c_{u}^{\varepsilon}\|\mu_{N}^{\varepsilon}(\hat{x}$$\|)$
@@
$\leq$ $2\|Q\|\sqrt{B^{\varepsilon}_{N}\frac{\bar{\lambda}}{\underline{\lambda}}}L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}\|\hat{x}\|^{2}+\underbrace{c_{u}^{\varepsilon}\|\hat{x}\|\|\mu_{N}^{\varepsilon}(\hat{x})\|}_{\leq\frac{1}{2}c_{u}^{\varepsilon}(\|\hat{x}\|^{2}+\|\mu_{N}^{\varepsilon}(\hat{x})\|^{2})})$
@@
$V_{N}^{\varepsilon}(f(\hat{x},\mu_{N}^{\varepsilon}(\hat{x})))\leq V_{N}^{\varepsilon}(\hat{x})-\alpha^{\varepsilon}\ell(\hat{x},\mu_{N}^{\varepsilon}(\hat{x}))+C_{x}\|\hat{x}\|^{2}+C_{u}\|\mu_{N}^{\varepsilon}(\hat{x})\|^{2},$
@@
$C_{x}:=2\|Q\|\sum_{i=0}^{N-1}\left(\sqrt{B^{\varepsilon}_{N}\frac{\bar{\lambda}}{\underline{\lambda}}}L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}+\frac{1}{2}c_{u}^{\varepsilon})+(L_{f^{\varepsilon}}^{i}c_{x}^{\varepsilon})^{2}\right),$
and
$C_{u}:=2\|Q\|\sum_{i=0}^{N-1}\left(\sqrt{B^{\varepsilon}_{N}\frac{\bar{\lambda}}{\underline{\lambda}}}L_{f^{\varepsilon}}^{i}\frac{1}{2}c_{u}^{\varepsilon}+(L_{f^{\varepsilon}}^{i}c_{u}^{\varepsilon})^{2}\right).$
$C_{x}$ and $C_{u}$ can be made arbitrarily small with sufficiently small proportionality constants $c_{x}^{\varepsilon}$ and $c_{u}^{\varepsilon}$. Then, there exists a sufficiently small $\varepsilon_{0}$ (potentially further tightened in comparison to our first choice resulting from the computation of the lower bound $\alpha_{N}^{\varepsilon_{0}}$ on the suboptimality index) such that $\bar{c}_{x}$, $\bar{c}_{u}$ are sufficiently small to ensure the inequality
@@
Proof. The proof resembles the proof of Proposition 6 in Müller and Worthmann (2017). In the chain of inequalities (12), see the proof of Theorem 1, we showed that there exists a sufficiently small $\varepsilon$ such that
$V^{\varepsilon}_{N}(f^{\varepsilon}(x,\mu^{\varepsilon}_{N}(x)))\leq V^{\varepsilon}_{N}(x)-\alpha^{\varepsilon}_{N}\ell(x,\mu^{\varepsilon}_{N}(x))$
holds for all $x\in V_{N}^{-\varepsilon}(c)$. Then, denoting by $\mathbf{u}_{\mathrm{MPC}}$ the input sequence obtained applying the MPC control law $\mu^{\varepsilon}_{N}$, we can relate the infinite-horizon optimal cost to $\alpha^{\varepsilon}_{N}$
@@
Hence, the growth bounds holds with $B_{k}^{\varepsilon}:=B_{N}^{\varepsilon}/\alpha^{\varepsilon}_{N}$, $k>N$, which shows the assertion. $\square$
@@
Finally, we emphasize the symmetry in the derived results meaning that verifying cost controllability of the surrogate model (3) for sufficiently small accuracy parameter $\varepsilon$ would suffice to conclude cost controllability of the original system and, thus, also to ensure asymptotic stability of the MPC closed loop based on the surrogate model, but applied to the original system (1).
@@ 4. Data-driven models via kernel EDMD
In this section, we verify Assumptions 1 and 2, i.e., the assumptions imposed on the surrogate model (3), using Koopman-operator theory. To this end, we leverage kernel extended dynamic mode decomposition (kEDMD) and its extension to control systems proposed in Bold et al. (2025b) as a data-driven method to approximate the Koopman operator. Moreover, we briefly discuss kernel approximations of the original dynamics (1) in Subsection 4.3 to emphasize that using the Koopman operator only serves as one of potentially many options to generate data-driven surrogate models satisfying Assumptions 1 and 2.
Let $\mathsf{k}:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a continuous, symmetric and strictly positive-definite kernel function, i.e., for every set $\mathcal{X}=\{x_{1},\dots,x_{d}\}\subset\mathbb{R}^{n}$ of pairwise distinct elements, the kernel matrix $K_{\mathcal{X}}=(\mathsf{k}(x_{i},x_{j}))_{i,j=1}^{p}$ is positive definite. For $x\in\mathbb{R}^{n}$, the canonical features $\Phi_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ of $\mathsf{k}$ are defined by $\Phi_{x}(x^{\prime})=\mathsf{k}(x,x^{\prime})$, $x^{\prime}\in\mathbb{R}^{n}$. By completion, the kernel $\mathsf{k}$ induces an Hilbert space $\mathbb{H}$ with inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$, see Wendland (2004). Importantly, elements $f\in\mathbb{H}$ fulfill the reproducing property
$f(x)=\langle f,\Phi_{x}\rangle_{\mathbb{H}}\qquad\forall\,x\in\mathbb{R}^{n}$
showing that point evaluation is well defined in $\mathbb{H}$. $\mathbb{H}$ is called reproducing kernel Hilbert space (RKHS).
In this work, we use piecewise-polynomial and compactly-supported kernel functions based on the Wendland radial basis functions (RBFs) $\Phi_{n,k}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ with smoothness degree $k\in\mathbb{N}$. The RKHS induced by the Wendland kernels coincides with fractional Sobolev spaces with equivalent norms Wendland (2004), a key property, which also holds for Matérn kernels; see Fasshauer and Ye (2011). The induced (Wendland) kernel is given by
$\mathsf{k}(x,y)=\phi_{n,k}(\|x-y\|)\qquad\text{for }x,y\in\mathbb{R}^{n}$
In the numerical simulations of Section 5, $k=1$ is used, which corresponds to $\phi_{n,1}\in\mathcal{C}^{2}([0,\infty),\mathbb{R})$ defined by
@@
4.1. Koopman operator and EDMD
@@
where $\mathcal{C}_{b}(\Omega)$ is the space of continuous and bounded real-valued functions on the set $\Omega$.
For a set of $d\in\mathbb{N}$ pairwise distinct points
@@
$V_{\mathcal{X}}:=\operatorname{span}\{\Phi_{x_{1}},\dots,\Phi_{x_{d}}\}\subset\mathbb{H}$ is the $d$-dimensional subspace of features. Further, $P_{\mathcal{X}}$ denotes the $\mathbb{H}$-orthogonal projection onto $V_{\mathcal{X}}$, i.e., for $h\in\mathbb{H}$, $P_{\mathcal{X}}h$ solves the regression problem $\min_{g\in V_{\mathcal{X}}}\|g-h\|^{2}_{\mathbb{H}}$. By the reproducing property, $P_{\mathcal{X}}h$ interpolates $h$ at $\mathcal{X}$.
Let $\mathcal{K}|_{V_{\mathcal{X}}}$ be the restriction of the Koopman operator on the subspace $V_{\mathcal{X}}$. As introduced in Köhne et al. (2025), we consider the matrix approximant $\widehat{K}$ of $\mathcal{K}|_{V_{\mathcal{X}}}$, which is given by
@@
We note that, for a large number of data points in $\mathcal{X}$, the computation of the approximation of the Koopman operator (17) can be numerically instable due to bad conditioning of the kernel matrix $K_{\mathcal{X}}$. To alleviate this, one may include regularization by replacing ${K}_{\mathcal{X}}^{-1}$ with $({K}_{\mathcal{X}}+\lambda I)^{-1}$ with regularization parameter $\lambda>0$. Error bounds in the spirit of Theorem 2 for this regularized surrogate were proven in (Bold et al., 2025b, Theorem 2.4).
4.2. Extension of kernel EDMD to control systems
In this section, we recap the extension of kEDMD as introduced in Subsection 4.1 to systems with inputs introduced in Bold et al. (2025b, c). The kernel-EDMD surrogate model is defined for control-affine discrete-time systems
@@ 4.2. Extension of kernel EDMD to control systems
with $\operatorname{rank}(U_{i})=m+1$ for all $i\in[1:d]$.
Assumption 3 requires $D:=\sum_{i=1}^{d}d_{i}$ data triplets. We call $x_{i}$ a virtual observation (or cluster) point to emphasize that no samples at $x_{i}$ are required. Further, $r_{\mathcal{X}}$ corresponds to the cluster radius, i.e., the radius of the neighborhoods around the virtual observation points, from which the data points are sampled.
@@
$H_{i}=[\tilde{g}_{0}(x_{i})\mid\tilde{G}(x_{i})]\in\mathbb{R}^{n\times(m+1)}$
of the matrix $g_{0}(x_{i})\mid G(x_{i})]$ for each $x_{i}\in\mathcal{X}$ by solving the linear regression problem
$\operatorname*{arg\,min}_{H_{i}}\big\|[x^{+}_{i1}\mid\ldots\mid x^{+}_{id_{i}}]-H_{i}U_{i}\big\|_{F}.$
The solution may be expressed by the pseudoinverse $U_{i}^{\dagger}$, which, for every virtual observation point $x_{i}$, is well defined in view of Assumption 3. Later, we will become interested in a uniform bound on $\max_{i\in[1:d]}\sqrt{d_{i}}\|U_{i}^{\dagger}\|$. On the one hand, this may be achieved by generating (sufficiently) exciting control values $u_{ij}$, $j\in[1:d_{i}]$, following the guidelines derived in Schmitz et al. (2025) leveraging, among others, subspace-angle conditions, to ensure data efficiency ($d_{i}=m+(1-\delta_{1i})$). On the other hand, one may also use randomly chosen input values in combination with a slightly larger $d_{i}$, see (Bold et al., 2025b, Remark 4.6).
To derive proportional error bounds for the surrogate model, we directly incorporate our knowledge for the data point $x_{1}=0$, that is, $g_{0}(x_{1})=0$. Hence, we set $\tilde{g}_{0}(x_{1})=0$ while $\tilde{G}(x_{1})$ is obtained by solving the reduced regression problem
@@
resembling the approach presented, e.g., in (Umlauft et al., 2018, Section III.B).
@@
$\leq$ $\Big[C_{J_{g_{0}}}+\max_{u\in\mathbb{U}}\ \sum_{i=1}^{m}C_{J_{g_{i}}}|u_{i}|\Big]h_{\mathcal{X}}^{k-1/2}+\|J_{\hat{g}_{0}}(z)-J_{g^{\varepsilon}_{0}}(z)\|+\sum\limits_{i=1}^{m}\|(J_{\hat{g}_{i}}(z)-J_{g^{\varepsilon}_{i}}(z))u_{i}\|,$
@@
$\leq$ $\max_{u\in\mathbb{U}}|u_{i}|L_{J_{g_{i}}}\sqrt{2\max_{k\in[1:d]}d_{k}}\left(\max_{\ell\in[1:d]}\|U^{\dagger}_{\ell}\|\right)\cdot\|K_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(z)\|\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}$
$\leq$ $\max_{u\in\mathbb{U}}|u_{i}|L_{J_{g_{i}}}\sqrt{2\max_{k\in[1:d]}d_{k}}\left(\max_{\ell\in[1:d]}\|U^{\dagger}_{\ell}\|\right)\Phi_{n,k}(0)\cdot\max_{v\in\mathds{1}}v^{\top}K_{\mathcal{X}}^{-1}v\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}=:c_{\mathcal{X}i},$
@@
$\widetilde{C}\!+\!C_{J_{g_{0}}}h_{\mathcal{X}}^{k-1/2}\!\!+m\!\!\!\max\limits_{i\in[1:m],u\in\mathbb{U}}C_{J_{g_{i}}}|u_{i}|h_{\mathcal{X}}^{k-1/2}\!\!+n\!\sum_{i=0}^{m}c_{\mathcal{X}i}.$
@@
$\leq\underbrace{C_{J_{g_{0}}}h_{\mathcal{X}}^{k-1/2}\!+\!m\max\limits_{i\in[1:m],u\in\mathbb{U}}C_{J_{g_{i}}}|u_{i}|h_{\mathcal{X}}^{k-1/2}+n\sum_{i=0}^{m}c_{\mathcal{X}i}.}_{=c_{x}^{\varepsilon}}$
To compute a bound for $\|J_{f^{\varepsilon},u}(x,u)-J_{f,u}(x,u)\|$, we exploit the error bounds of Theorem 3, to obtain
@@
4.3. Relation to kernel approximation of control systems
@@
for each $i\in[1:n]$. This corresponds to direct interpolation of the flow map and hence involves only one projection. Moreover, the main difference to the approximation (17) are the data requirements: For the approximation (17) applied to the coordinate functions, we require data through the lens of the features, i.e., $\mathsf{k}(x_{i},F(x_{j}))$ for $(i,j)\in[1:d]^{2}$, whereas for the approximation (26), we require samples of the flow map, i.e., $F(x_{j})$ for all $j\in[1:d]$. The error bound for the surrogate dynamics (26), however, is very closely related to the one inferred from (17) and of similar structure, see (Köhne et al., 2025, Theorem 3.4), where the first estimate in this reference corresponds to (26) (involving only one projection) and the second estimate to (17) (involving two projections). As both projection errors are proportional to the fill distance $h_{\mathcal{X}}$, the convergence rate is same and the results of Subsection 4.2 can be derived analogously. Hence, Assumption 1 can also be verified for the surrogate model (26).

### v3

@@ 3. Asymptotic stability of MPC
Since, we have the inequality $\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+}),u^{\sharp}(i))\geq\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|_{Q}^{2}$ for all $i\in[0:N-1]$, we also have $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|_{Q}^{2}\leq B^{\varepsilon}_{N}\|\hat{x}\|_{Q}^{2}$. By exploiting standard inequalities of weighted squared norms and taking the square root, we have $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|\leq\sqrt{B^{\varepsilon}_{N}\bar{\lambda}/\underline{\lambda}}\|\hat{x}\|$. Then, substituting this inequalities in (15), we get
@@
$\leq$ $)\displaystyle 2\|Q\|\sqrt{B^{\varepsilon}_{N}\bar{\lambda}/\underline{\lambda}}\|\hat{x}\|L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}\|\hat{x}\|+c_{u}^{\varepsilon}\|\mu_{N}^{\varepsilon}(\hat{x}$$\|)$
@@
$\leq$ $2\|Q\|\sqrt{B^{\varepsilon}_{N}\bar{\lambda}/\underline{\lambda}}L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}\|\hat{x}\|^{2}+\underbrace{c_{u}^{\varepsilon}\|\hat{x}\|\|\mu_{N}^{\varepsilon}(\hat{x})\|}_{\leq\frac{1}{2}c_{u}^{\varepsilon}(\|\hat{x}\|^{2}+\|\mu_{N}^{\varepsilon}(\hat{x})\|^{2})})$
@@
[…]
$\leq$ $V_{N}^{\varepsilon}(\hat{x})-\alpha^{\varepsilon}\ell(\hat{x},\mu_{N}^{\varepsilon}(\hat{x}))+C_{x}\|\hat{x}\|^{2}+C_{u}\|\mu_{N}^{\varepsilon}(\hat{x})\|^{2},$
@@
$C_{x}:=2\|Q\|\sum_{i=0}^{N-1}\left(\sqrt{B^{\varepsilon}_{N}\bar{\lambda}/\underline{\lambda}}L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}+\frac{1}{2}c_{u}^{\varepsilon})+(L_{f^{\varepsilon}}^{i}c_{x}^{\varepsilon})^{2}\right)$
and $C_{u}:=2\|Q\|\cdot\sum_{i=0}^{N-1}\big(\sqrt{B^{\varepsilon}_{N}\frac{\bar{\lambda}}{\underline{\lambda}}}L_{f^{\varepsilon}}^{i}\frac{1}{2}c_{u}^{\varepsilon}+(L_{f^{\varepsilon}}^{i}c_{u}^{\varepsilon})^{2}\big)$. $C_{x}$ and $C_{u}$ can be made arbitrarily small with sufficiently small proportionality constants $c_{x}^{\varepsilon}$ and $c_{u}^{\varepsilon}$. Then, there exists a sufficiently small $\varepsilon_{0}$ (potentially further tightened in comparison to our first choice resulting from the computation of the lower bound $\alpha_{N}^{\varepsilon_{0}}$ on the suboptimality index) such that $\bar{c}_{x}$, $\bar{c}_{u}$ are sufficiently small to ensure the inequality
@@
Proof. The proof resembles the proof of Proposition 6 in Müller and Worthmann (2017). In the chain of inequalities (12), see the proof of Theorem 1, we showed that there exists a sufficiently small $\varepsilon$ such that $V^{\varepsilon}_{N}(f^{\varepsilon}(x,\mu^{\varepsilon}_{N}(x)))\leq V^{\varepsilon}_{N}(x)-\alpha^{\varepsilon}_{N}\ell(x,\mu^{\varepsilon}_{N}(x))$ holds for all $x\in V_{N}^{-\varepsilon}(c)$. Then, denoting by $\mathbf{u}_{\mathrm{MPC}}$ the input sequence obtained applying the MPC control law $\mu^{\varepsilon}_{N}$, we can relate the infinite-horizon optimal cost to $\alpha^{\varepsilon}_{N}$
@@
Hence, the growth bound holds with $B_{k}^{\varepsilon}:=B_{N}^{\varepsilon}/\alpha^{\varepsilon}_{N}$ for all $k>N$, which shows the assertion. $\square$
@@
We emphasize the symmetry in the derived results meaning that verifying cost controllability of the surrogate model (3) for sufficient accuracy parameter $\varepsilon$ would suffice to conclude cost controllability of the original system and, thus, also to ensure asymptotic stability of the MPC closed loop based on the surrogate model, but applied to the original system (1).
@@ 4. Data-driven models via kernel EDMD
In this section, we verify Assumptions 1 and 2, i.e., the assumptions imposed on the surrogate model (3), using Koopman-operator theory. To this end, we leverage kernel extended dynamic mode decomposition (kEDMD) and its extension to control systems proposed in Bold et al. (2025b) as a data-driven method to approximate the Koopman operator. Moreover, at the end of the section we briefly discuss kernel approximations of the original dynamics (1) to emphasize that using the Koopman operator only serves as one of potentially many options to generate data-driven surrogate models satisfying Assumptions 1 and 2.
Let $\mathsf{k}:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a continuous, symmetric and strictly positive-definite kernel function, i.e., for every set $\mathcal{X}=\{x_{1},\dots,x_{d}\}\subset\mathbb{R}^{n}$ of pairwise distinct elements, the kernel matrix $K_{\mathcal{X}}=(\mathsf{k}(x_{i},x_{j}))_{i,j=1}^{p}$ is positive definite. For $x\in\mathbb{R}^{n}$, the canonical features $\Phi_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ of $\mathsf{k}$ are defined by $\Phi_{x}(x^{\prime})=\mathsf{k}(x,x^{\prime})$, $x^{\prime}\in\mathbb{R}^{n}$. By completion, the kernel $\mathsf{k}$ induces an Hilbert space $\mathbb{H}$ with inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$, see Wendland (2004). Importantly, elements $f\in\mathbb{H}$ fulfill the reproducing property $f(x)=\langle f,\Phi_{x}\rangle_{\mathbb{H}}$ for all $x\in\mathbb{R}^{n}$ showing that point evaluation is well defined in $\mathbb{H}$. $\mathbb{H}$ is called reproducing kernel Hilbert space (RKHS). In this work, we use piecewise-polynomial and compactly-supported kernel functions based on the Wendland radial basis functions (RBFs) $\Phi_{n,k}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ with smoothness degree $k\in\mathbb{N}$. The RKHS induced by the Wendland kernels coincides with fractional Sobolev spaces with equivalent norms Wendland (2004), a key property, which also holds for Matérn kernels; see Fasshauer and Ye (2011). The induced (Wendland) kernel is given by $\mathsf{k}(x,y)=\phi_{n,k}(\|x-y\|)$ for $x,y\in\mathbb{R}^{n}$. In the numerical simulations of Section 5, $k=1$ is used, which corresponds to $\phi_{n,1}\in\mathcal{C}^{2}([0,\infty),\mathbb{R})$ defined by
@@
Koopman operator and EDMD
@@
where $\mathcal{C}_{b}(\Omega)$ is the space of continuous and bounded functions on the set $\Omega$. For a set of $d\in\mathbb{N}$ pairwise distinct points
@@
$V_{\mathcal{X}}:=\operatorname{span}\{\Phi_{x_{1}},\dots,\Phi_{x_{d}}\}\subset\mathbb{H}$ is the $d$-dimensional subspace of features. Further, $P_{\mathcal{X}}$ denotes the $\mathbb{H}$-orthogonal projection onto $V_{\mathcal{X}}$, i.e., for $h\in\mathbb{H}$, $P_{\mathcal{X}}h$ solves the regression problem $\min_{g\in V_{\mathcal{X}}}\|g-h\|^{2}_{\mathbb{H}}$. By the reproducing property, $P_{\mathcal{X}}h$ interpolates $h$ at $\mathcal{X}$. Let $\mathcal{K}|_{V_{\mathcal{X}}}$ be the restriction of the Koopman operator on the subspace $V_{\mathcal{X}}$. As introduced in Köhne et al. (2025), we consider the matrix approximant $\widehat{K}$ of $\mathcal{K}|_{V_{\mathcal{X}}}$, which is given by
@@
We note that, for large data sets $\mathcal{X}$, the computation of the approximation of the Koopman operator (17) can be numerically instable due to bad conditioning of the kernel matrix $K_{\mathcal{X}}$. To alleviate this, one may include regularization by replacing ${K}_{\mathcal{X}}^{-1}$ with $({K}_{\mathcal{X}}+\lambda I)^{-1}$ with regularization parameter $\lambda>0$. Error bounds in the spirit of Theorem 2 for this regularized surrogate were proven in (Bold et al., 2025b, Theorem 2.4).
Extension of kernel EDMD to control systems.
In this section, we recap the extension of kEDMD to systems with inputs introduced in Bold et al. (2025b, c). The kernel-EDMD surrogate model is defined for control-affine discrete-time systems
@@
satisfy $\operatorname{rank}(U_{i})=m+1$ for all $i\in[1:d]$.
Assumption 3 requires $D:=\sum_{i=1}^{d}d_{i}$ data triplets. We call $x_{i}$ a virtual observation (or cluster) point to emphasize that no samples at $x_{i}$ are assumed. Further, $r_{\mathcal{X}}$ corresponds to the cluster radius of the neighborhoods around the virtual observation points $x_{i}$ from which the data are sampled.
@@
To derive proportional error bounds for the surrogate model, we incorporate our knowledge for the data point $x_{1}=0$, that is, $g_{0}(x_{1})=0$. Hence, we set $\tilde{g}_{0}(x_{1})=0$ while $\tilde{G}(x_{1})$ is obtained by solving the reduced regression problem
@@
resembling the approach presented in (Umlauft et al., 2018, Section III.B).
@@
$\leq$ $\Big[C_{J_{g_{0}}}+\max_{u\in\mathbb{U}}\ \sum_{i=1}^{m}C_{J_{g_{i}}}|u_{i}|\Big]h_{\mathcal{X}}^{k-1/2}$
$+\|J_{\hat{g}_{0}}(z)-J_{g^{\varepsilon}_{0}}(z)\|+\sum\limits_{i=1}^{m}\|(J_{\hat{g}_{i}}(z)-J_{g^{\varepsilon}_{i}}(z))u_{i}\|,$
@@
$\leq$ $\max_{u\in\mathbb{U}}|u_{i}|L_{J_{g_{i}}}\sqrt{2\max_{k\in[1:d]}d_{k}}\left(\max_{\ell\in[1:d]}\|U^{\dagger}_{\ell}\|\right)$
$\cdot\|K_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(z)\|\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}$
$\leq$ $\max_{u\in\mathbb{U}}|u_{i}|L_{J_{g_{i}}}\sqrt{2\max_{k\in[1:d]}d_{k}}\left(\max_{\ell\in[1:d]}\|U^{\dagger}_{\ell}\|\right)\Phi_{n,k}(0)$
$\cdot\max_{v\in\mathds{1}}v^{\top}K_{\mathcal{X}}^{-1}v\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}=:c_{\mathcal{X}i},$
@@
$\widetilde{C}+C_{J_{g_{0}}}h_{\mathcal{X}}^{k-1/2}+m\!\max\limits_{i\in[1:m],u\in\mathbb{U}}C_{J_{g_{i}}}|u_{i}|h_{\mathcal{X}}^{k-1/2}+n\sum\nolimits_{i=0}^{m}c_{\mathcal{X}i}.$
@@
$\leq C_{J_{g_{0}}}h_{\mathcal{X}}^{k-1/2}+m\max\limits_{i\in[1:m],u\in\mathbb{U}}C_{J_{g_{i}}}|u_{i}|h_{\mathcal{X}}^{k-1/2}+n\sum_{i=0}^{m}c_{\mathcal{X}i},$
i.e., $c_{x}^{\varepsilon}$. To compute a bound for $\|J_{f^{\varepsilon},u}(x,u)-J_{f,u}(x,u)\|$, we exploit the error bounds of Theorem 3, to obtain
@@
Relation to kernel approximation of control systems.
@@
for each $i\in[1:n]$. This corresponds to direct interpolation of the flow map and hence involves only one projection. Moreover, the main difference to the approximation (17) are the data requirements: For the approximation (17) applied to the coordinate functions, we require data through the lens of the features, i.e., $\mathsf{k}(x_{i},F(x_{j}))$ for $(i,j)\in[1:d]^{2}$, whereas for the approximation (26), we require samples of the flow map, i.e., $F(x_{j})$ for all $j\in[1:d]$. The error bound for the surrogate dynamics (26), however, is very closely related to the one inferred from (17) and of similar structure, see (Köhne et al., 2025, Theorem 3.4), where the first estimate in this reference corresponds to (26) (involving only one projection) and the second estimate to (17) (involving two projections). As both projection errors are proportional to the fill distance $h_{\mathcal{X}}$, the convergence rate is same and the results above can be derived analogously. Hence, Assumption 1 can also be verified for the surrogate model (26).